# An ALMA spectroscopic survey of the Planck high-redshift object PLCK G073.457.5 confirms two protoclusters

### v1

@@ 3.2 ALMA Band 4 continuum search
We also searched our primary beam-uncorrected data cubes for previously undiscovered (i.e., not reported in Kneissl et al. 2019) submm continuum sources. To do this, we calculated the noise-weighted mean of each pointing (LSB+USB), where the noise per channel was estimated as the standard deviation of all the pixels after masking out the 18 known galaxies from Kneissl et al. (2019), and obviously bright galaxies in our SCUBA-2 pointings. We searched each continuum map for both positive and negative S/N peaks, finding that the most significant negative peak across all twelve maps was 5.1. Following our line search, we set this as our continuum search threshold. This search did not yield any new detections in the pointings centred on the positions of the eight previous ALMA observations (fields 1 through 8), but in the four new pointings centred on SCUBA-2 sources, we detected five sources (one in fields S2-1, S2-2, and S2-3, and two in S2-0).
@@ 3.3 Redshifts, line measurements and continuum measurements
First, for each position in our catalogue, we create 3′′′′continuum-subtracted cutouts in our Band 4 primary beam-corrected data cubes using the CASA task imcontsub. For the continuum subtraction, at each pixel we fit a 0th-order polynomial (i.e., a constant) across frequency space to the channels containing no line emission, and subtract the fit.
@@
Next, we generate LSB and USB spectra for each galaxy using the continuum-subtracted data cubes by manually placing elliptical apertures at the positions returned by either our line search or our continuum search, and summing the pixels within the aperture. The radius and ellipticity of the apertures are set by the size of the 2 contours in each cutout after calculating the average over all available channels. For sources 4 (located in field 2) and 24 (located in field S2-3), two significant emission lines were detected in the spectrum. For both cases, averaging over the channels corresponding to each line, showed that the two moment-0 maps were spatially offset by about 0.5′′, with one peak spatially coincident with the bright continuum source and the other having no continuum counterpart. We checked that there were no pairs of millimetre lines able to explain the two peaks assuming they came from the same galaxy (e.g. CO(7–6) and [Cii](2–1) around ). We therefore classify the two secondary sources as companion galaxies (designated 4b and 24b), bringing the total number of line detections to 13. The resulting spectra are provided in Appendix A.
@@
As discussed above, we assume that there are two possible CO transitions for the lines detected in the LSB, while the lines in the USB are Ci(1–0) at . To establish which CO transition we have observed, we note that for lines detected between about 137.8 and 139.1 GHz we would expect CO(4–3) in the LSB and Ci(1–0) in the USB; this is the case for five sources (IDs 8, 9, 12, 13, and 18). For these sources, while Ci(1–0) was not detected in the USB with S/N, lines are still possible at a lower significance. We note that CO(4–3) typically has a peak flux density brighter by a factor of 2 compared to Ci(1–0) in SMGs (e.g., Birkin et al., 2021; Hagimoto et al., 2023), while for each of these sources the line detected in the LSB has a peak flux density about 3–5 times the spectrum rms.
@@
We therefore perform a likelihood ratio test. We first fit a single Gaussian with three free parameters (amplitude, mean, and standard deviation) to the LSB line and calculate the likelihood function (here given by ). We then perform a second fit for two Gaussians, where only the amplitude of the second Gaussian is a free parameter as the mean is fixed to the expected frequency of Ci(1–0) and the standard deviation is fixed to be equal to the first Gaussian. We recalculate the likelihood function, and then calculate . The difference in degrees of freedom between the models is 1, so is expected to follow a distribution with 1 degree of freedom and the relevant statistic () is the survival function (or 1 minus the cumulative distribution function) evaluated at . This statistic can be interpreted as the probability that the null hypothesis (in this case the model with one Gaussian) describes the data better than the alternative, and so we apply a threshold of 0.05 below which we reject the null hypothesis.
@@
We find that we cannot reject the null hypothesis for IDs 8, 9, and 12 (, , and , respectively), thus we interpret the line emission as CO(3–2). For IDs 13 and 18 we reject the null hypothesis with 95% confidence ( and , respectively), so we assume that we have detected both CO(3–2) and Ci(1–0). Kneissl et al. (2019) discussed a tentative (S/N) CO(5–4) line detection in ID 8 at , but the CO(3–2) line we have identified with an S/N of about 17 places this galaxy at – we conclude that the possible CO(5–4) line was in fact a noise excursion.
@@
For IDs 4 and 24 (plus their companion galaxies) the line emission is observed at a frequency where we cannot perform this test. ID 4 (and its companion 4b) happens to lie behind a large foreground galaxy and therefore we do not have good near-IR continuum measurements, thus we just assume that it is at the same redshift as ID 3. For ID 24 we fit the available near-IR and mm continuum photometry to a range of possible redshifts assuming the detected line is CO(3–2), CO(4–3), and Ci(1–0), then calculate for each fit. We find that a redshift of 2.6, corresponding to the Ci(1–0) line, minimizes , so we adopt this as the redshift of this galaxy and its companion. Lastly, the line detected at the location of ID 3 is unambiguously CO(3–2), since two other CO lines have previously been detected.
@@
For the 13 galaxies with line emission, we fit single Gaussian functions to the observed lines in order to estimate various line properties. From the fit, line strengths/luminosities are estimated by summing all channels within of the mean of the fit, while the line FWHM are taken directly from the standard deviation of the fit, and redshifts are calculated using the mean of each fit. The line luminosities are calculated in units following Solomon et al. (1997) as
@@
where is the line intensity derived from the Gaussian fit, the observed frequency of the line (see Table 1) and is the usual cosmological luminosity distance. We also calculate the line luminosities in solar luminosity units following the same integration range.
@@ 4.1 Stellar masses and star-formation rates
We estimate the stellar mass of the line-detected sources by modelling their SEDs with the Code Investigating GAlaxy Emission (CIGALE; Boquien et al., 2019). This code models galaxy SEDs using stellar and dust components, and offers the advantage of modelling simultaneously the NIR part of the SED and the sub-mm/mm emission in a self-consistent way by preserving the energy balance between the stellar radiation that is absorbed by the dust and its re-emission at longer wavelengths (see Buat et al. 2019 for a discussion on this assumption).
@@
For the reference model, we assumed a delayed star-formation history (SFH) with an optional burst, a Chabrier (2003) initial mass function (IMF), the stellar population models of Bruzual & Charlot (2003), the dust models from Draine et al. (2014), the Calzetti et al. (2000) attenuation law, and solar metallicity ( = 0.02). The best-fit is assessed through , and the best-fit parameters and associated uncertainties are estimated as the likelihood-weighted means and standard deviations. The CIGALE best-fit models, available for seven out of nine CO-emitters with available NIR photometry, are shown in Fig. 2. The stellar masses and SFRs (instantaneous, SFRinst, and averaged over the past 100 Myrs, SFR100) derived from the CIGALE best-fits are reported in Table 3. We also derive SFRIR, the SFR from the total IR luminosity () assuming the relation in Kennicutt & Evans (2012) corrected for a Chabrier IMF (i.e., ; see Table 3). The IR luminosity () was obtained from fitting the sub-mm/mm data with a modified blackbody using the cmcirsed package (Casey, 2012), and assuming the CO-derived redshift and a dust emissivity-index equal to 1.8 (Cortese et al., 2014; Pokhrel et al., 2016). Since SFRinst and SFR100 can differ significantly, by up to a factor of 3, we use SFRIR in the following because it is usually intermediate between the two values provided by the CIGALE best fit. In addition, it is directly comparable with the values reported in the literature where it is commonly used, in particular for mm-selected galaxies. Compared to SFRIR, SFRinst can be twice as high, and SFR100 four times smaller. We note that Kneissl et al. (2019) estimated the SFRs for the same galaxy sample using the same cmcirsed package; however, they lacked spectroscopic redshifts and the additional 2-mm flux densities (). Nonetheless, we find comparable results.
@@
In Fig. 3 we show the offset from the average star-forming main-sequence (MS) as parameterized by Popesso et al. (2023) using each source’s redshift and stellar mass. The grey shaded region corresponds to the scatter of 0.3 dex around the MS. Following Rodighiero et al. (2011), we classify a source as a starburst if the SFR is four times higher than expected according to the MS for a galaxy with the same stellar mass and at the same redshift, and as a normal SFG if the SFR is lower than this threshold. Four G073 sources (IDs 1, 8, 12, and 24) are considered starbursts based on the MS offset and the remaining six are consistent with the MS. In the figure, we also show the position of CO-detected cluster and protocluster members at from the literature (see the caption for detailed references). Compared to other clusters at similar redshifts, the G073 sources are on average more star-forming, but sources with similarly high SFRs are observed in other structures.
@@ 4.2 Molecular gas masses
where is the scaling parameter. To derive gas masses from Ci we use a similar equation, namely
We note that since we have not observed the Ci(2–1) transition we cannot take into account the gas excitation temperature, which is commonly used to calculate the partition function and obtain gas mass estimates with less uncertainty (e.g., Bothwell et al., 2017; Dunne et al., 2021; Gururajan et al., 2023). This simple parameterization was calibrated using a large sample of SMGs with both Ci(1–0) and CO(4–3) detections by Birkin et al. (2021), finding an average of 5.21.3, which we use here.
With this calibration, the only conversion factors we need to assume are and the scaling of the various CO luminosities to . can range from around 0.2 to 10 M⊙ pc-2 (K km s-1)-1 (Tacconi et al., 2008; Casey et al., 2014), depending on the physical properties of the galaxy such as metallicity (Genzel et al., 2012, 2015; Amorín et al., 2016; Tacconi et al., 2018; Inoue et al., 2021) or SFR, with respect to that expected for a galaxy with the same stellar mass and redshift on the main sequence (MS) relation (see equation 2 in Castignani et al., 2020). Galaxies above the MS, like starburst galaxies, can have smaller values than those on the MS. Similarly, galaxies with lower metallicities can have smaller values. The –metallicity relation is, however, valid only for massive (1010 M⊙) SFGs (Genzel et al., 2015). For the galaxies in G073 with 1010 M⊙ the derived values range from 4.0 and 6.7 (Genzel et al., 2015; Tacconi et al., 2018). To be consistent with previous works, we adopt a conservative = 4.36 M⊙ pc-2 (K km s-1)-1 (Bolatto et al., 2013; Genzel et al., 2015) for all sources, independent of their MS-based classification. This conversion factor is commonly used for the Milky Way and for normal SFGs with solar metallicities (Bolatto et al., 2013; Genzel et al., 2015), and it includes a correction factor for helium. Regardless of the choice of , there are large systematic uncertainties involved in estimating the gas mass, which we ignore here for simplicity.
@@ 5.1 Redshift distribution
In Fig. 4 we show the distribution of the 13 galaxy redshifts we have detected in our ALMA survey, alongside the photometric redshift probability distributions originally determined for the 18 galaxies from Kneissl et al. (2019). The two structures initially predicted by Kneissl et al. (2019) at and are seen as the two broad peaks in the probability distribution, and the redshifts we have spectroscopically confirmed agree with these two peaks. For the structure, we find six galaxies with an average redshift of , while for the structure we find five galaxies (ignoring the SCUBA-2 galaxies) with an average redshift of . Evidently none of the SCUBA-2 sources we have targeted are within the structures identified by Planck and Herschel, which we discuss further below.
Figure 5 shows the spatial distribution of the two structures, colour-coded according to their line-of-sight velocity relative to the mean redshift of each group. We find that most of the galaxies are spatially concentrated in the north-west corner of the field-of-view initially defined by the Planck beam, while the galaxies are concentrated in the south-east corner, about 1′from the lower-redshift structure. This distribution points to the conclusion that the original Planck data has selected an area on the sky where two protoclusters are aligned along the line of sight to within about 3′of one-another, relative to the 5′Planck beam.

### v2

@@ 3.2 ALMA Band 4 continuum search
We also searched our primary beam-uncorrected data cubes for previously undiscovered (i.e., not reported in Kneissl et al. 2019) submm continuum sources. To do this, we calculated the noise-weighted mean of each pointing (LSB+USB), where the noise per channel was estimated as the standard deviation of all the pixels after masking out the 18 known galaxies from Kneissl et al. (2019), and obviously bright galaxies in our SCUBA-2 pointings. We searched each continuum map for both positive and negative S/N peaks, finding that the most significant negative peak across all 12 maps was 5.1. Following our line search, we set this as our continuum search threshold. This search did not yield any new detections in the pointings centred on the positions of the eight previous ALMA observations (fields 1 through 8), but in the four new pointings centred on SCUBA-2 sources, we detected five sources (one in fields S2-1, S2-2, and S2-3, and two in S2-0).
@@ 3.3 Redshifts, line measurements and continuum measurements
First, for each position in our catalogue, we created 3′′′′continuum-subtracted cutouts in our Band 4 primary beam-corrected data cubes using the CASA task imcontsub. For the continuum subtraction, at each pixel we fit a zeroth-order polynomial (i.e. a constant) across frequency space to the channels containing no line emission, and subtracted the fit.
@@
Next, we generated LSB and USB spectra for each galaxy using the continuum-subtracted data cubes by manually placing elliptical apertures at the positions returned by either our line search or our continuum search, and summing the pixels within the aperture. The radius and ellipticity of the apertures are set by the size of the 2 contours in each cutout after calculating the average over all available channels. For reference, the average aperture semi-major axis size was found to be 1.6′′. For sources 4 (located in field 2) and 24 (located in field S2-3), two significant emission lines were detected in the spectrum. For both cases, averaging over the channels corresponding to each line, showed that the two moment-0 maps were spatially offset by about 0.5′′, with one peak spatially coincident with the bright continuum source and the other having no continuum counterpart. We checked that there were no pairs of millimetre lines able to explain the two peaks assuming they came from the same galaxy (e.g. CO(7–6) and [Cii](2–1) around ). We therefore classify the two secondary sources as companion galaxies (designated 4b and 24b), bringing the total number of line detections to 13. The resulting spectra are provided in Appendix A.
@@
As discussed above, we assumed that there are two possible CO transitions for the lines detected in the LSB, while the lines in the USB are Ci(1–0) at . To establish which CO transition we observed, we note that for lines detected between about 137.8 and 139.1 GHz we would expect CO(4–3) in the LSB and Ci(1–0) in the USB; this is the case for five sources (IDs 8, 9, 12, 13, and 18). For these sources, while Ci(1–0) was not detected in the USB with S/N, lines are still possible at a lower significance. We note that CO(4–3) typically has a peak flux density brighter by a factor of 2 compared to Ci(1–0) in SMGs (e.g., Birkin et al. 2021; Hagimoto et al. 2023), while for each of these sources the line detected in the LSB has a peak flux density about 3–5 times the spectrum rms.
@@
We therefore performed a likelihood ratio test. We first fit a single Gaussian with three free parameters (amplitude, mean, and standard deviation) to the LSB line and calculate the likelihood function (here given by ). We then performed a second fit for two Gaussians, where only the amplitude of the second Gaussian is a free parameter as the mean is fixed to the expected frequency of Ci(1–0) and the standard deviation is fixed to be equal to the first Gaussian. We recalculated the likelihood function and then calculated . The difference in degrees of freedom between the models is 1, so is expected to follow a distribution with 1 degree of freedom and the relevant statistic () is the survival function (or 1 minus the cumulative distribution function) evaluated at . This statistic can be interpreted as the probability that the null hypothesis (in this case the model with one Gaussian) describes the data better than the alternative, and so we applied a threshold of 0.05 below which we rejected the null hypothesis.
@@
We find that we cannot reject the null hypothesis for IDs 8, 9, and 12 (, , and , respectively); thus, we interpret the line emission as CO(3–2). For IDs 13 and 18 we reject the null hypothesis with 95% confidence ( and , respectively), so we assumed that we have detected both CO(3–2) and Ci(1–0). Kneissl et al. (2019) discussed a tentative (S/N) CO(5–4) line detection in ID 8 at , but the CO(3–2) line we have identified with an S/N of about 17 places this galaxy at – we conclude that the possible CO(5–4) line was in fact a noise excursion. Comparing our other results to Kneissl et al. (2019): the spectroscopic redshift of ID 1 is 2.306, compared to the photometric redshift of 2.420.15; 1.506 compared to 2.210.22 for ID 9; 2.306 compared to 2.430.29 for ID 11; 1.507 compared to 1.400.10 for ID 12; and 2.340 compared to 2.630.25 for ID 13. Finally, due to the revised NIR and MIR counterparts for ID 4, the photometric redshift was not reliable and cannot be compared to our spectroscopic redshift.
@@
For IDs 4 and 24 (plus their companion galaxies) the line emission is observed at a frequency where we cannot perform this test. ID 4 (and its companion 4b) happens to lie behind a large foreground galaxy and therefore we do not have good NIR continuum measurements; thus, we just assumed that it is at the same redshift as ID 3. For ID 24 we fit the available NIR and mm continuum photometry to a range of possible redshifts assuming the detected line is CO(3–2), CO(4–3), and Ci(1–0), then calculate for each fit. We find that a redshift of 2.6, corresponding to the Ci(1–0) line, minimizes , so we adopted this as the redshift of this galaxy and its companion. Lastly, the line detected at the location of ID 3 is unambiguously CO(3–2), since two other CO lines have previously been detected.
@@
For the 13 galaxies with line emission, we fit single Gaussian functions to the observed lines in order to estimate various line properties. From the fit, line strengths are estimated by summing all channels within of the mean of the fit, while the line FWHM are taken directly from the standard deviation of the fit, and redshifts are calculated using the mean of each fit. The line luminosities are calculated in units following Solomon et al. (1997) as
@@
where is the line intensity derived from the Gaussian fit, the observed frequency of the line (see Table 1) and is the usual cosmological luminosity distance. We also calculated the line luminosities in solar luminosity units following the same integration range.
@@ 4.1 Stellar masses and star-formation rates
We estimated the stellar mass of the line-detected sources by modelling their SEDs with the Code Investigating GAlaxy Emission (CIGALE; Boquien et al. 2019). This code models galaxy SEDs using stellar and dust components, and offers the advantage of modelling simultaneously the NIR part of the SED and the submm/millimetre emission in a self-consistent way by preserving the energy balance between the stellar radiation that is absorbed by the dust and its re-emission at longer wavelengths (see Buat et al. 2019 for a discussion on this assumption).
@@
For the reference model, we assumed a delayed star-formation history with an optional burst, a Chabrier (2003) IMF, the stellar population models of Bruzual & Charlot (2003), the dust models from Draine et al. (2014), the Calzetti et al. (2000) attenuation law, and solar metallicity ( = 0.02). The best-fit is assessed through , and the best-fit parameters and associated uncertainties are estimated as the likelihood-weighted means and standard deviations. The CIGALE best-fit models, available for seven out of nine CO-emitters with available NIR photometry, are shown in Fig. 2. The stellar masses and SFRs (instantaneous, SFRinst, and averaged over the past 100 Myrs, SFR100) derived from the CIGALE best-fits are reported in Table 3.
We also derived SFRIR, the SFR from the total infrared luminosity () assuming the relation in Kennicutt & Evans (2012) corrected for a Chabrier IMF (i.e. ; see Table 3). The infrared luminosity was obtained from fitting the submm/millimetre data with a modified blackbody using the cmcirsed package (Casey 2012), and assuming the CO-derived redshift and a dust emissivity-index equal to 1.8 (Cortese et al. 2014; Pokhrel et al. 2016), leaving the dust temperatures as free parameters. The resulting average dust temperature of the sample was found to be (273) K. Since SFRinst and SFR100 can differ significantly, by up to a factor of 3, we use SFRIR in the following because it is usually intermediate between the two values provided by the CIGALE best fit. In addition, it is directly comparable with the values reported in the literature where it is commonly used, in particular for millimetre-selected galaxies. Compared to SFRIR, SFRinst can be twice as high, and SFR100 four times smaller. We note that Kneissl et al. (2019) estimated the SFRs for the same galaxy sample using the same cmcirsed package; however, they lacked spectroscopic redshifts and the additional 2-mm flux densities (). Nonetheless, we find comparable results. We also tested running CIGALE with the cmcirsed dust model, which includes shorter wavelength data. This resulted in similar or slightly higher SFRs, owing to the fact that the model includes a power-law component in the MIR that is not used when fitting the FIR/millimetre data alone; however, the differences were within the uncertainties.
@@
In Fig. 3 we show the offset from the average star-forming main sequence (MS) as parameterized by Popesso et al. (2023) using each source’s redshift and stellar mass. The grey shaded region corresponds to the scatter of 0.3 dex around the MS. Following Rodighiero et al. (2011), we classified a source as a starburst if the SFR is four times higher than expected according to the MS for a galaxy with the same stellar mass and at the same redshift, and as a normal star-forming galaxy (SFG) if the SFR is lower than this threshold. Four G073 sources (IDs 1, 8, 12, and 24) are considered starbursts based on the MS offset and the remaining six are consistent with the MS. In the figure we also show the position of CO-detected cluster and protocluster members at from the literature (see the caption for detailed references). Compared to other clusters at similar redshifts, the G073 sources are on average more star-forming, but sources with similarly high SFRs are observed in other structures.
@@ 4.2 Molecular gas masses
where is the scaling parameter. To derive gas masses from Ci we used a similar equation, namely
We note that since we have not observed the Ci(2–1) transition we cannot take into account the gas excitation temperature, which is commonly used to calculate the partition function and obtain gas mass estimates with less uncertainty (e.g., Bothwell et al. 2017; Dunne et al. 2021; Gururajan et al. 2023). This simple parameterization was calibrated using a large sample of SMGs with both Ci(1–0) and CO(4–3) detections by Birkin et al. (2021), finding an average of 5.21.3, which we used here.
With this calibration, the only conversion factors we need to assume are and the scaling of the various CO luminosities to . can range from around 0.2 to 10 M⊙ pc-2 (K km s-1)-1 (Tacconi et al. 2008; Casey et al. 2014), depending on the physical properties of the galaxy such as metallicity (Genzel et al. 2012, 2015; Amor´ın et al. 2016; Tacconi et al. 2018; Inoue et al. 2021) or SFR, with respect to that expected for a galaxy with the same stellar mass and redshift on the MS relation (see equation 2 in Castignani et al. 2020). Galaxies above the MS, like starburst galaxies, can have smaller values than those on the MS. Similarly, galaxies with lower metallicities can have smaller values. The –metallicity relation, however, is valid only for massive (1010 M⊙) SFGs (Genzel et al. 2015). For the galaxies in G073 with 1010 M⊙ the derived values range from 4.0 and 6.7 (Genzel et al. 2015; Tacconi et al. 2018). To be consistent with previous works, we adopted a conservative = 4.36 M⊙ pc-2 (K km s-1)-1 (Bolatto et al. 2013; Genzel et al. 2015) for all sources, independent of their MS-based classification. This conversion factor is commonly used for the Milky Way and for normal SFGs with solar metallicities (Bolatto et al. 2013; Genzel et al. 2015), and it includes a correction factor for helium. Regardless of the choice of , there are large systematic uncertainties involved in estimating the gas mass, which we ignore here for simplicity.
As a last step, the measured CO luminosities were converted to CO(1–0) luminosities, , using the and to ratios and (Birkin et al. 2021). The derived gas masses are listed in Table 3.
For comparison, we also tested standard scaling relations between the Rayleigh-Jeans dust continuum luminosity and gas mass. Since we only have dust continuum estimates for our entire sample from the ALMA Band 6 observations as 1.3 mm (rest frame 390–520 m; see Kneissl et al. 2019), we scaled these to rest-frame 850 m following Equations 9 and 10 from Liu et al. (2019) assuming a modified blackbody SED with and a dust temperature of 35 K, then apply the dust continuum-to-gas mass scaling factor of [erg s-1 Hz-1 M] from Scoville et al. (2016). We find similar results, with a mean ratio of the dust continuum gas mass over the CO/Ci gas mass of 1.04. For the following comparisons with literature values we used our CO/Ci-derived gas masses since they are not correlated with our SFR estimates, which rely significantly on the same dust continuum flux densities; however our results are not strongly influenced by our choice of gas mass estimator.
@@ 5.1 Redshift distribution
In Fig. 4 we show the distribution of the 13 galaxy redshifts we detected in our ALMA survey, alongside the combined photometric redshift probability distributions originally determined for all of the 18 galaxies from Kneissl et al. (2019). Here the combined probability density was calculated by summing the probability densities estimated by EAZY over the whole sample. The two structures initially predicted by Kneissl et al. (2019) at and are seen as the two broad peaks in the probability distribution. Since this distribution includes the photometric redshift uncertainties, the widths of the peaks correspond to the uncertainties in the redshifts of the protocluster structures – roughly 0.2 for the structure, and 0.5 for the structure. The redshifts we have spectroscopically confirmed agree with these two peaks. For the structure, we find six galaxies with an average redshift of , while for the structure we find five galaxies (ignoring the SCUBA-2 galaxies) with an average redshift of . Evidently none of the SCUBA-2 sources we targeted are within the structures identified by Planck and Herschel, which we discuss further below.
Figure 5 shows the spatial distribution of the two structures, colour-coded according to their line-of-sight velocity relative to the mean redshift of each group. We find that most of the galaxies are spatially concentrated in the north-west corner of the field of view initially defined by the Planck beam, while the galaxies are concentrated in the south-east corner, about 1′from the lower-redshift structure. This distribution points to the conclusion that the original Planck data has selected regions on the sky where two protoclusters are aligned along the line of sight to within about 3′of one another, relative to the 5′Planck beam.